August 7, 2017. Logbook system is on the fritz and the imaging system on the scope isn’t connecting. I’ll get it all straight tomorrow
August 5, 2017. Joe got the logbook system working last night. Let me know if you have any more trouble.
August 4, 2017. Reports of logbook system not working right. We are going to make a fix tonight I’ll let you know if it works. Otherwise since I’m out of town it might be Monday
July 31, 2017. We are conducting some vibration isolations system test this afternoon and tomorrow morning so the TEM will be unavailable. It should be back online tomorrow afternoon.
July 24, 2017. Please use the reservation calendar no matter when you plan to use the instruments. I use it to track use in the calendar system. I’ve told a few of you and it continues to happen. This is not optional.
July 24, 2017. 2:09 pm FIB is back online (NM) July 24, 2017. 11:08 am There is an air leak on theFIB. I am working on it now.
July 13, 2017. The FE-SEM is now back online and working normally.
July 13, 2017. Right now the FE-SEM is having issues with the stage control and will be offline until further notice.
June 15, 2017. After 2 hours on the phone, the Hitachi engineer and I found out why the data bar was positioned so far into the images. Someone had changed a software setting – a box was checked that shouldn’t have been. Needless to say, he and I had other …
June 5, 2017. If you make any changes to the basic operational configuration of the FESEM software please return it to the original settings. Two things that seem to happen too frequently. One is changing the detector from Upper to Lower or Mixed. It should be left on Upper. The …